# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations via Projected Gradient Descent (PGD), and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and VQA tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks. Our code and models can be found at: https://github.com/luisrui/Modality-Interference-in-MLLMs.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have made significant strides in integrating vision and language understanding within a unified architecture  llava ; instructblip ; qwen25vl . By combining powerful visual encoders and large language models through alignment mechanisms, MLLMs such as LLaVA llava and Qwen-VL qwen25vl demonstrate strong capabilities across a wide range of multimodal tasks. However, beneath their seemingly impressive performance lies a critical limitation: MLLMs often fail to distinguish between relevant and irrelevant signals across modalities, leading to unreliable predictions wang2024mdpo ; vlmconflict ; spurlens . Moreover, while MLLMs are designed for multimodal tasks, their failure on unimodal tasks—where only a single modality (e.g. text) should guide the prediction—raises concerns about whether the model can preserve modality-specific competencies. For instance, MLLMs frequently underperform on pure visual recognition imagewikiqa ; mof and textual reasoning vlmconflict ; cogvlm ; vila , suggesting that cross-modal fusion may induce unintended interference and degrade unimodal performance.
@@
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting imagewikiqa ; mof ; cogvlm ; vila , knowledge conflict wang2024mdpo ; vlmconflict , and spurious correlations more ; spurlens . Catastrophic forgetting has been identified as a key factor in visual degradation imagewikiqa ; mof , where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict wang2024mdpo ; vlmconflict impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations more ; spurlens show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA llava fail to fully align vision and language representations, resulting in unstable modality reliance mof ; vlmconflict ; EECA . Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding imagewikiqa and diminished language understanding vila . Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency Competency —the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
@@
Figure 1: Perturbation results on image classification (left) and text QA (right) tasks. Left: Mini-ImageNet with Orig (Original input), UF (Add Unrelated Facts), and MD (Add Misleading Descriptions). Right: OpenBookQA with RP (Random Pixels), IR (Attach Irrelevant Real Image), BK (Attach Black Canvas), and WH (Attach White Canvas). For image-heavy tasks, spurious textual cues (esp. misleading descriptions) severely degrade accuracy. For text-heavy tasks, unrelated real images disrupt reasoning—highlighting modality interference in MLLMs.
@@
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles more to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure˜1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images significantly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
@@
The empirical results from Figure˜1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentations strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations includes both heuristic perturbations (e.g. add unrelated facts in image-heavy prompts) and adversarial training time perturbations, which exposes the model to worst-case alignment disruptions, serving as a strong variant. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families llava15 ; instructblip ; qwen25vl and diverse benchmarks miniimagenet ; caltech101 ; openbookqa ; mmlu ; scienceqa ; mmbench ; seedbench demonstrate the superiority of our method.
@@ 2 Related Works
Recent studies have revealed that modality misalignment remains a key obstacle in MLLMs, leading to degraded performance on both image-heavy and text-heavy tasks. For visual understanding, catastrophic forgetting occurs when multimodal tuning overrides pretrained visual features imagewikiqa ; mof ; cogvlm . In text-heavy scenarios, knowledge conflict vlmconflict arises when inconsistent parametric knowledge from different modalities confuse reasoning. Moreover, mDPO wang2024mdpo identifies language bias in training, where models fail to condition their responses on visual input. Some works attribute such issues to shallow fusion cogvlm —e.g., LLaVA llava uses lightweight projectors to bridge vision and language spaces, leaving a representational gap and resulting in loosely coupled features mof ; vlmconflict . Others highlight data limitations: even well-encoded visual features fail to support reasoning without adequate supervision to guide decoding imagewikiqa .
Building on these diagnoses, recent models have proposed multiple solutions. MoF mof mitigates this by fusing features from multiple vision encoders, while VLMClassifier imagewikiqa enhances recognition via vision-only finetuning, though it struggles with VQA due to lack of cross-modal alignment. CogVLM cogvlm introduces a visual expert module to improve vision-language integration. VILA vila , QwenVL qwen25vl , and InternVL internvl3 incorporate text-only supervision in different ways to preserve or enhance language capabilities during multimodal training—through stage-wise separation, parallel preservation, and unified joint optimization, respectively. These works motivate us to hypothesize on the root cause—the model’s inability to assess modality relevance. We further propose a causal framing of modality interference and introduce a perturbation-based fine-tuning strategy to improve modality discriminability and inference-time robustness.
@@
Adversarial perturbations threaten the reliability of both vision and text tasks by exposing vulnerabilities through deliberate and imperceptible perturbations. Existing robustness methods can be broadly categorized by modality, targeting either continuous image embeddings or discrete token spaces.
In vision tasks, attacks like FGSM DBLP:journals/corr/GoodfellowSS14 and CW DBLP:conf/sp/Carlini017 first revealed the fragility of neural networks to imperceptible input changes. PGD DBLP:conf/iclr/MadryMSTV18 formalized this under a saddle-point framework, becoming the standard for adversarial training. AutoAttack DBLP:conf/icml/Croce020a further unified strong attacks, including PGD variants, into a reliable benchmark. In text tasks, adversarial methods must contend with discrete inputs. TextFooler DBLP:conf/aaai/JinJZS20 substitutes key words with semantically similar ones to mislead predictions, while CodeAttack DBLP:conf/aaai/JhaR23 adapts this idea to code-language models. More recently, PGD has been extended to LLMs via continuous relaxation DBLP:journals/corr/abs-2402-09154 , enabling efficient, gradient-based attacks in embedding space. Beyond evaluation, PGD has also been used as a regularizer to improve optimization. PTP DBLP:conf/emnlp/ChenCHC23 applies PGD-style perturbations in the prompt embedding space to smooth training and enhance stability. Inspired by this, our work extends PGD to the multimodal embedding space, enabling unified gradient-based control over both visual and textual inputs.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features Competency . We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency as an ability for MLLM which no single modality should be sufficient to determine the correct prediction, and models are expected to integrate information across modalities to make robust decisions. A competent model should be able to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section˜5 and Section˜C.1.
@@
We formalize modality interference through a causal intervention perspective with a causal graph, as shown in Figure˜2, where visual inputs ($X_{I}$) and textual inputs ($X_{T}$) are processed into their respective representations ($Z_{I}$, $Z_{T}$) before being fused to produce the final prediction ($A$). To study the model’s reliance on different modalities, we introduce perturbations directly at the input level, serving as causal interventions causalIntervention on $X_{I}$ and $X_{T}$. Under ideal cross-modality competency, the model’s prediction should primarily depend on the task-relevant pathway (e.g., $X_{I}\rightarrow Z_{I}\rightarrow A$ in image-heavy tasks, $X_{T}\rightarrow Z_{T}\rightarrow A$ in text-heavy tasks). Causal interventions at the input level allow us to diagnose whether the model improperly fuses irrelevant signals into its decision process. We use $x_{I}\textquoteright$ to denote an intervention on image $X_{I}$ and use $x_{T}\textquoteright$ as an intervention on text $X_{T}$. Following Pearl’s causal framework causalIntervention ; intermediaryVariable , we quantify the impact of modality perturbations on model predictions by formalizing causal effects in our multimodal setting. Specifically, we define the pre-intervention prediction distribution as $P(A|X_{I},X_{T})$, and the post-intervention prediction distribution after applying a perturbation on $X_{I}$ or $X_{T}$ as $P\textquoteright(A|\text{do}(X_{I}=x_{I}\textquoteright)$ or $\text{do}(X_{T}=x_{T}\textquoteright))$. The do-operation represents an intervention to specific modality, and the causal effect (CE) of an intervention is evaluated via a distance metric $\delta$ comparing $P$ and $P\textquoteright$ as $\text{CE}=\delta(P,P^{\prime})$. We assess the causal effect via prediction changes using $\delta_{\text{cp}}(P,P^{\prime}):=\mathbb{I}\left(a\neq a^{\prime}\right)$ in which $a=\arg\max_{x}P(x)$ is the predicted answer before intervention, $a\textquoteright=\arg\max_{x}P\textquoteright(x)$ is the predicted answer after intervention and $\mathbb{I}(\cdot)$ is the indicator function that outputs $1$ if $a\neq a\textquoteright$ and 0 otherwise. Thus, $\delta_{\text{cp}}$ captures whether the model’s final decision $A$ changes under perturbations to the input modality. In all interventions, a high value of $\delta_{\text{cp}}$ indicates the model’s susceptibility to modality interference, revealing spurious reliance on irrelevant modality.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations, and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and multimodal tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks. Our code and models can be found at: https://github.com/luisrui/Modality-Interference-in-MLLMs
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have made significant strides in integrating vision and language understanding within a unified architecture  (Liu et al., 2023b; Luo et al., 2023; Bai et al., 2025). By combining powerful visual encoders and large language models through alignment mechanisms, MLLMs such as LLaVA (Liu et al., 2023b) and Qwen-VL (Bai et al., 2025) demonstrate strong capabilities across a wide range of multimodal tasks. However, beneath their seemingly impressive performance lies a critical limitation: MLLMs often fail to distinguish between relevant and irrelevant signals across modalities, leading to unreliable predictions (Wang et al., 2024a; Zhu et al., 2024; Hosseini et al., 2025). Moreover, while MLLMs are designed for multimodal tasks, their failure on unimodal tasks—where only a single modality (e.g. text) should guide the prediction—raises concerns about whether the model can preserve modality-specific competencies. For instance, MLLMs frequently underperform on pure visual recognition (Zhang et al., 2024; Tong et al., 2024) and textual reasoning (Zhu et al., 2024; Wang et al., 2023; Lin et al., 2024), suggesting that cross-modal fusion may induce unintended interference and degrade unimodal performance.
@@
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023; Lin et al., 2024), knowledge conflict (Wang et al., 2024a; Zhu et al., 2024), and spurious correlations (Chen et al., 2024a; Hosseini et al., 2025). Catastrophic forgetting has been identified as a key factor in visual degradation (Zhang et al., 2024; Tong et al., 2024), where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict (Wang et al., 2024a; Zhu et al., 2024) impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations (Chen et al., 2024a; Hosseini et al., 2025; Zhou et al., 2025) show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA (Liu et al., 2023b) fail to fully align vision and language representations, resulting in unstable modality reliance (Tong et al., 2024; Zhu et al., 2024; Zhao et al., 2025). Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding (Zhang et al., 2024) and diminished language understanding (Lin et al., 2024). Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency (Gardner et al., 2021)—the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
@@
Figure 1: Performance degradation under irrelevant perturbations reveals modality interference in MLLMs. Left: Mini-ImageNet (image-heavy) with Original input, Unrelated Facts, and Misleading Descriptions. Right: OpenBookQA (text-heavy) with Random Pixels, Full Black Canvas, and Irrelevant Real Images. Misleading descriptions induces the most severe degradation in image-heavy tasks, while irrelevant real images cause the largest drop in text-heavy reasoning.
@@
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles (Pearl, 1995; Chen et al., 2024a) to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure 1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images mostly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
@@
The empirical results from Figure 1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentation strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations include both heuristic variants (e.g., injecting unrelated facts into image-heavy prompts) and adversarial training-time perturbations, which expose the model to worst-case alignment disruptions and thus serve as a stronger form of regularization. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families and diverse benchmarks demonstrate the superiority of our method.
@@ 2 Related Works
Recent studies have revealed that modality misalignment remains a key obstacle in MLLMs, leading to degraded performance on both image-heavy and text-heavy tasks. For visual understanding, catastrophic forgetting occurs when multimodal tuning overrides pretrained visual features (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023). In text-heavy scenarios, knowledge conflict (Zhu et al., 2024) arises when inconsistent parametric knowledge from different modalities confuse reasoning. Moreover, mDPO (Wang et al., 2024a) identifies language bias in training, where models fail to condition their responses on visual input. Some works attribute such issues to shallow fusion (Wang et al., 2023)—e.g., LLaVA uses lightweight projectors to bridge vision and language spaces, leaving a representational gap and resulting in loosely coupled features (Tong et al., 2024; Zhu et al., 2024; Zhao et al., 2025). Others highlight data limitations: even well-encoded visual features fail to support reasoning without adequate supervision to guide decoding (Zhang et al., 2024).
Building on these diagnoses, recent models have proposed multiple solutions. MoF (Tong et al., 2024) mitigates this by fusing features from multiple vision encoders, while VLMClassifier (Zhang et al., 2024) enhances recognition via vision-only finetuning, though it struggles with VQA due to lack of cross-modal alignment. CogVLM (Wang et al., 2023) introduces a visual expert module to improve vision-language integration. VILA (Lin et al., 2024), QwenVL (Bai et al., 2025), and InternVL (Zhu et al., 2025) incorporate text-only supervision in different ways to preserve or enhance language capabilities during multimodal training—through stage-wise separation, parallel preservation, and unified joint optimization, respectively. Similar patterns arise in multimodal structural reasoning, where models must rely on relevant modalities to generalize to unseen relations (Cai et al., 2024). These works motivate us to hypothesize on the root cause—the model’s inability to assess modality relevance. We further propose a causal framing of modality interference and introduce a perturbation-based fine-tuning strategy to improve the inference-time robustness of MLLMs.
@@
Adversarial perturbations threaten the reliability of both vision and text tasks by exposing vulnerabilities through deliberate and imperceptible perturbations. Existing robustness methods can be broadly categorized by modality, targeting either continuous image embeddings or discrete token spaces. In vision tasks, attacks like FGSM (Goodfellow et al., 2015) and CW (Carlini & Wagner, 2017) first revealed the fragility of neural networks to imperceptible input changes. PGD (Madry et al., 2018) formalized this under a saddle-point framework, becoming the standard for adversarial training. AutoAttack (Croce & Hein, 2020) further unified strong attacks, including PGD variants, into a reliable benchmark. In text tasks, adversarial methods must contend with discrete inputs. TextFooler (Jin et al., 2020) substitutes key words with semantically similar ones to mislead predictions, while CodeAttack (Jha & Reddy, 2023) adapts this idea to code-language models. More recently, PGD has been extended to LLMs via continuous relaxation (Geisler et al., 2024), enabling efficient attacks in embedding space. Beyond evaluation, PGD has also been used as a regularizer to improve optimization. PTP (Chen et al., 2023a) applies PGD-style perturbations in the prompt embedding space to smooth training and enhance stability. Inspired by this, our work extends PGD to the multimodal embedding space, enabling unified gradient-based control over both visual and textual inputs.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features (Gardner et al., 2021). We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency as an ability for MLLM to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section 5, Section C.1 and Table 8.
@@
[…]